OpenStudy (anonymous):

I know what is the product of 11/15 and 12/7 is 44/35, I just need to understand why?

OpenStudy (anonymous):

the product is 132/105, which reduces to 44/35

OpenStudy (anonymous):

@omg--do i just cross multiply and get the 132/105? then reduce down to 44/35?

OpenStudy (vishweshshrimali5):

see \[\frac{a}{b} * \frac{c}{d} = \frac{a*c}{b*d} = \frac{ac}{bd}\] Therefore, \[\frac{11}{15} * \frac{12}{7} = \frac{11*12}{15*7} = \frac{11 * 4*3}{5*3*7} = \frac{11*4}{5*7} = \frac{44}{35}\]

OpenStudy (anonymous):

take the two numerators and multiply them. 11*12 = 132 same with the denominators. 15*7 = 105 then you reduce to 44/35.

OpenStudy (anonymous):

yeppers and that is how it's done, you were on the right track al along...:)

OpenStudy (anonymous):

i see--multiply them straight across, not diagonally.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

thanks you guys!

OpenStudy (vishweshshrimali5):

did u get it ?

OpenStudy (anonymous):

welcome

OpenStudy (vishweshshrimali5):

our pleasure